# Probabilistic Pontryagin’s Maximum Principle for Continuous-Time Model-Based Reinforcement Learning

### v2

@@ III-B Probabilistic Hamiltonian
The posterior distribution over system dynamics $p(\boldsymbol{\theta}\mid\mathcal{D})$ in Eq. (II.6) implicitly defines a distribution over Hamiltonian functions $p(\mathcal{H}\mid\mathcal{D})$ through its dependence on the dynamics. Sampling a realization from the dynamics posterior yields a corresponding Hamiltonian:
$\mathcal{H}(\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{u},t;\boldsymbol{\theta})\coloneq L\left(\boldsymbol{x},\boldsymbol{u}\right)+\boldsymbol{\lambda}^{\top}\boldsymbol{f}\left(\boldsymbol{x},\boldsymbol{u},t;\boldsymbol{\theta}\right).$ (III.5)
@@
Our goal is to find a control function $\boldsymbol{u}^{*}(t)$ that minimizes the mean cost functional $\bar{C}(\boldsymbol{u})$. For a single realization $\boldsymbol{\theta}$, the necessary optimality conditions for $C(\boldsymbol{u};\boldsymbol{\theta})$ follow from the minimum principle:
$\boldsymbol{u}^{*}(t;\boldsymbol{\theta})=\operatorname*{argmin}_{\boldsymbol{u}}\mathcal{H}\left(\boldsymbol{x}^{*}(t;\boldsymbol{\theta}),\boldsymbol{\lambda}^{*}(t;\boldsymbol{\theta}),\boldsymbol{u},t;\boldsymbol{\theta}\right)$
@@
with corresponding state and costate trajectories $\boldsymbol{x}^{*}(t;\boldsymbol{\theta})$ and $\boldsymbol{\lambda}^{*}(t;\boldsymbol{\theta})$ satisfying Eqs. (III.1) – (III.3). For the necessary optimality conditions of $\bar{C}(\boldsymbol{u})$, we consider the mean minimum principle:
@@
$\boldsymbol{u}^{*}(t)=\operatorname*{argmin}_{\boldsymbol{u}}\mathbb{E}_{\theta}\left[\mathcal{H}\left(\boldsymbol{x}(t;\boldsymbol{\theta}),\boldsymbol{\lambda}(t;\boldsymbol{\theta}),\boldsymbol{u},t;\boldsymbol{\theta}\right)\right]$ (III.6)
where $\boldsymbol{x}(t;\boldsymbol{\theta})$ and $\boldsymbol{\lambda}(t;\boldsymbol{\theta})$ are the state and costate trajectories that satisfy Eqs. (III.1) – (III.3) for a realization $\boldsymbol{\theta}\sim p(\boldsymbol{\theta}\mid\mathcal{D})$. These optimality conditions result in a family of state and co-state trajectories, that are optimal in expectation but not for each individual realization of $\boldsymbol{\theta}$. Despite this, the co-states $\boldsymbol{\lambda}(t;\boldsymbol{\theta})$ remain interpretable as cost sensitivity by satisfying the transversality condition.
III-C Necessary optimality conditions for the mean cost
We now consider if the mean minimum principle is a necessary condition for optimality of the mean cost $\bar{C}(\boldsymbol{u})$. We take a variational approach that assumes unbounded controls, but note that the constrained setting can be approached using the needle (or spike) variation of the proof [16].
Theorem 1 (Mean Pontryagin Minimum Principle).
Let $\boldsymbol{u}^{*}(t)$ be an optimal control that minimizes the mean cost functional $\bar{C}(\boldsymbol{u})$, and let $\mathcal{U}=\mathbb{R}^{d_{u}}$. Then a necessary condition for optimality is:
@@ III-C Necessary optimality conditions for the mean cost
We assume that $\mathcal{H}$ is differentiable following $\boldsymbol{f}$ and $L$’s differentiability, and that $\nabla_{\boldsymbol{u}}\mathcal{H}$ is integrable w.r.t. $p(\boldsymbol{\theta}\mid\mathcal{D})$. We consider the Gateaux derivative of the mean cost:
$\delta\bar{C}(\boldsymbol{u})=\lim\limits_{\epsilon\to 0}\frac{\bar{C}(\boldsymbol{u}+\epsilon\delta\boldsymbol{u})-\bar{C}(\boldsymbol{u})}{\epsilon}$
where $\delta\boldsymbol{u}(t)$ is an arbitrary variation defined over $[t_{0},t_{f}]$. The first variation can be obtained by isolating the first order terms around the optimal solution $\boldsymbol{u}^{*}(t)$:
$\bar{C}(\boldsymbol{u}^{*}+\epsilon\delta\boldsymbol{u})-\bar{C}(\boldsymbol{u}^{*})\approx\delta\bar{C}\big{|}_{\boldsymbol{u}^{*}(t)}(\delta\boldsymbol{u})\epsilon\enspace.$
Since by definition $\boldsymbol{u}^{*}(t)$ is optimal for $\bar{C}$, its first variation must be zero.
We first consider the deterministic case for a realization $\boldsymbol{\theta}$, where $\boldsymbol{u}^{*}(t;\boldsymbol{\theta})$ is optimal for $C(\boldsymbol{u};\boldsymbol{\theta})$. A well-known result in optimal control theory states that the first variation is given as [3, Eq. (5.1.13)]:
$\delta C_{\boldsymbol{\theta}}\big{|}_{\boldsymbol{u}^{*}(t;\boldsymbol{\theta})}(\delta\boldsymbol{u})=\int^{t_{f}}_{t_{0}}\bigl{[}$ $\nabla_{\boldsymbol{u}}\mathcal{H}_{*}\bigr{]}\delta\boldsymbol{u}(t)\operatorname{d}\!{t}\enspace,$
where $\mathcal{H}_{*}=\mathcal{H}\bigl{(}\boldsymbol{x}^{*}(t;\boldsymbol{\theta}),\boldsymbol{\lambda}^{*}(t;\boldsymbol{\theta}),\boldsymbol{u}^{*}(t;\boldsymbol{\theta}),t;\boldsymbol{\theta}\bigr{)}$ and $C_{\boldsymbol{\theta}}=C(\boldsymbol{u};\boldsymbol{\theta})$. By the fundamental lemma of calculus of variations, it follows that:
$\nabla_{\boldsymbol{u}}\mathcal{H}_{*}=0,\quad\text{for }t\in[t_{0},t_{f}]\enspace.$
Using this result, we now turn to the mean cost. Under standard regularity conditions (via dominated convergence), we interchange the derivative and expectation:
$\delta\bar{C}(\boldsymbol{u})=\delta\mathbb{E}_{\boldsymbol{\theta}}[C(\boldsymbol{u};\boldsymbol{\theta})]=\mathbb{E}_{\boldsymbol{\theta}}[\delta C(\boldsymbol{u};\boldsymbol{\theta})]\enspace.$
@@
$=\mathbb{E}_{\boldsymbol{\theta}}\bigl{[}\int^{t_{f}}_{t_{0}}\nabla_{\boldsymbol{u}}\mathcal{H}\bigl{(}\boldsymbol{x}(t;\boldsymbol{\theta}),\boldsymbol{\lambda}(t;\boldsymbol{\theta}),\boldsymbol{u}^{*}(t),t;\boldsymbol{\theta}\bigr{)}\delta\boldsymbol{u}\operatorname{d}\!{t}\bigr{]}$
$=\int^{t_{f}}_{t_{0}}\nabla_{\boldsymbol{u}}\mathbb{E}_{\boldsymbol{\theta}}\bigl{[}\mathcal{H}\bigl{(}\boldsymbol{x}(t;\boldsymbol{\theta}),\boldsymbol{\lambda}(t;\boldsymbol{\theta}),\boldsymbol{u}^{*}(t),t;\boldsymbol{\theta}\bigr{)}\bigr{]}\delta\boldsymbol{u}\operatorname{d}\!{t}$
@@
using the assumption that $\nabla_{\boldsymbol{u}}\mathcal{H}$ is integrable w.r.t. $p(\boldsymbol{\theta}\mid\mathcal{D})$. By the fundamental lemma of the calculus of variations, it follows that:
@@
This shows that the optimal control $\boldsymbol{u}^{*}(t)$ is only required to satisfy the optimality conditions in the expectation of $p(\boldsymbol{\theta}\mid\mathcal{D})$, extending the minimum principle to probabilistic dynamical models that capture epistemic uncertainty.
@@ IV-A Trajectory sampling
Computing the mean over Hamiltonians $\bar{\mathcal{H}}$ in Eq. (III.6) requires estimating the uncertainty in state and costate trajectories $\boldsymbol{x}^{*}(t;\boldsymbol{\theta})$ and $\boldsymbol{\lambda}^{*}(t;\boldsymbol{\theta})$ over $t\in[t_{0},t_{f}]$. We adopt the trajectory sampling technique to capture this variability introduced by the dynamics posterior [12]. The mean Hamiltonian is approximated by aggregating the sampled Hamiltonian functions:
$\boldsymbol{u}^{*}(t)\approx\operatorname*{argmin}_{\boldsymbol{u}}\frac{1}{M}\sum_{i=1}^{M}\mathcal{H}\left(\boldsymbol{x}(t;\boldsymbol{\theta}_{i}),\boldsymbol{\lambda}(t;\boldsymbol{\theta}_{i}),\boldsymbol{u},t;\boldsymbol{\theta}_{i}\right)$ (IV.1)
where $\boldsymbol{\theta}_{i}\sim p(\boldsymbol{\theta}\mid\mathcal{D})$ for $i=1,\dots,M$, and each conditioned Hamiltonian is computed along the trajectories $\boldsymbol{x}(t;\boldsymbol{\theta}_{i})$ and $\boldsymbol{\lambda}(t;\boldsymbol{\theta}_{i})$, which comply with the PMP conditions. An alternative formulation is to consider the mean of the optimal controls $\bar{\boldsymbol{u}}^{*}(t)=\mathbb{E}_{\boldsymbol{\theta}}[\boldsymbol{u}^{*}(t;\boldsymbol{\theta})]$:
$\bar{\boldsymbol{u}}^{*}(t)\approx\frac{1}{M}\sum^{M}_{i=1}\boldsymbol{u}^{*}(t;\boldsymbol{\theta}_{i})\enspace.$ (IV.2)
By not evaluating the mean Hamiltonian with a shared optimal control input for the mean cost, the mean optimal control approximation is only a valid necessary optimality condition for the mean cost when the Hamiltonian function is convex. By uncoupling the optimization process between different realizations of the dynamics posterior, however, the solutions can be computed in parallel, therefore leading to improved computational complexity. Both approaches leverage Monte Carlo estimates of the state and co-state trajectories to efficiently evaluate the $M$ samples using GPU parallelization.

### v3

@@ III-B Probabilistic Hamiltonian
In the presence of uncertainty, the posterior distribution over system dynamics $p(\boldsymbol{\theta}\mid\mathcal{D})$ in (II.6) implicitly defines a distribution over Hamiltonian functions $p(\mathcal{H}\mid\mathcal{D})$ through its dependence on the dynamics. Sampling a realization $\boldsymbol{\theta}\sim p(\boldsymbol{\theta}\mid\mathcal{D})$ yields a corresponding Hamiltonian:
$\mathcal{H}_{\boldsymbol{\theta}}(\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{u},t)\coloneq L\left(\boldsymbol{x},\boldsymbol{u}\right)+\boldsymbol{\lambda}^{\top}\boldsymbol{f}_{\boldsymbol{\theta}}\left(\boldsymbol{x},\boldsymbol{u},t\right).$ (III.5)
@@
which extends the classical Hamiltonian by incorporating the parameter-dependent dynamics $\boldsymbol{f}_{\boldsymbol{\theta}}$. Our goal is to find a control function $\boldsymbol{u}^{*}(t)$ that minimizes the mean cost functional $\mkern 1.5mu\overline{\mkern-1.5muC\mkern-1.5mu}\mkern 1.5mu(\boldsymbol{u})$. For a single realization $\boldsymbol{\theta}$, the state and co-state equations are:
$\dot{\boldsymbol{x}}_{\boldsymbol{\theta}}(t)$ $=\boldsymbol{f}_{\boldsymbol{\theta}}\bigl{(}\boldsymbol{x}_{\boldsymbol{\theta}}(t),\boldsymbol{u}(t),t\bigr{)},$ (III.6)
$\dot{\boldsymbol{\lambda}}_{\boldsymbol{\theta}}(t)$ $=-\nabla_{\boldsymbol{x}}\mathcal{H}_{\boldsymbol{\theta}}\bigl{(}\boldsymbol{x}_{\boldsymbol{\theta}}(t),\boldsymbol{\lambda}_{\boldsymbol{\theta}}(t),\boldsymbol{u}(t),t\bigr{)},$ (III.7)
$\boldsymbol{\lambda}_{\boldsymbol{\theta}}(t_{f})$ $=\nabla\Phi\bigl{(}\boldsymbol{x}_{\boldsymbol{\theta}}(t_{f})\bigr{)}.$ (III.8)
@@
Here, $\boldsymbol{x}_{\boldsymbol{\theta}}(t)$ and $\boldsymbol{\lambda}_{\boldsymbol{\theta}}(t)$ are the state and costate trajectories for a give admissible control $\boldsymbol{u}(t)$ and realization $\boldsymbol{\theta}$ satisfying (III.6) – (III.8). The necessary optimality condition for a single realization $C_{\boldsymbol{\theta}}(\boldsymbol{u})$ follows from the minimum principle:
@@
$\boldsymbol{u}_{\boldsymbol{\theta}}^{*}(t)\coloneq\operatorname*{argmin}_{\boldsymbol{u}}\mathcal{H}_{\boldsymbol{\theta}}\left(\boldsymbol{x}_{\boldsymbol{\theta}}^{*}(t),\boldsymbol{\lambda}_{\boldsymbol{\theta}}^{*}(t),\boldsymbol{u},t\right),$
for $t\in[t_{0},t_{f}]$, with $\boldsymbol{x}_{\boldsymbol{\theta}}^{*}(t)$ and $\boldsymbol{\lambda}_{\boldsymbol{\theta}}^{*}(t)$ satisfying (III.6) – (III.8). However, optimizing for a single $\boldsymbol{\theta}$ does not address the mean cost $\mkern 1.5mu\overline{\mkern-1.5muC\mkern-1.5mu}\mkern 1.5mu(\boldsymbol{u})$, which requires a control function that is optimal across the distribution of $\boldsymbol{\theta}$. To this end, we consider the mean minimum principle:
$\boldsymbol{u}^{*}(t)=\operatorname*{argmin}_{\boldsymbol{u}}\mathbb{E}_{\boldsymbol{\theta}}\left[\mathcal{H}_{\boldsymbol{\theta}}\left(\boldsymbol{x}_{\boldsymbol{\theta}}(t),\boldsymbol{\lambda}_{\boldsymbol{\theta}}(t),\boldsymbol{u},t\right)\right],$ (III.9)
This condition yields a control function that is optimal in expectation, producing a family of state and co-state trajectories $\boldsymbol{x}_{\boldsymbol{\theta}}(t)$ and $\boldsymbol{\lambda}_{\boldsymbol{\theta}}(t)$ that satisfy (III.6) – (III.8) for each $\boldsymbol{\theta}$. The mean minimum principle balances performance across all possible dynamics models weighted by their posterior probability, which is illustrated in Figure 2.
III-C Necessary optimality condition for the mean cost
We now state the necessary condition for optimality for the mean cost $\mkern 1.5mu\overline{\mkern-1.5muC\mkern-1.5mu}\mkern 1.5mu(\boldsymbol{u})$. The general principle developed in  [10] and [11] requires the optimal control $\boldsymbol{u}^{*}(t)$ to minimize the mean Hamiltonian at each point in time over the (possibly bounded) set of admissible controls $\mathcal{U}$. Here, we restrict ourselves to the case of unconstrained control setting, in which case this simplifies to a first-order condition on the gradient using a variational approach [20].
Assumptions: we assume that $\mathcal{U}=\mathbb{R}^{d_{u}}$, $\boldsymbol{f}_{\boldsymbol{\theta}}$ and $L$ are continuously differentiable w.r.t. $\boldsymbol{x}$ and $\boldsymbol{u}$, ensuring $\mathcal{H}_{\boldsymbol{\theta}}$ is differentiable for all $\boldsymbol{\theta}\in\Theta$, and $\nabla_{\boldsymbol{u}}\mathcal{H}_{\boldsymbol{\theta}}$ satisfies standard regularity conditions for interchanging $\mathbb{E}_{\boldsymbol{\theta}}$ with differentiation and integration over $t\in[t_{0},t_{f}]$.
Theorem 1 (Necessary Mean Cost Optimality Condition).
@@ III-C Necessary optimality condition for the mean cost
Since $\boldsymbol{u}^{*}(t)$ minimizes the mean cost, $\boldsymbol{x}_{\boldsymbol{\theta}}(t)$ and $\lambda_{\boldsymbol{\theta}}(t)$ are not necessarily optimal for each $\boldsymbol{\theta}$.
@@
We consider the Gateaux derivative of the mean cost:
$\delta\mkern 1.5mu\overline{\mkern-1.5muC\mkern-1.5mu}\mkern 1.5mu(\boldsymbol{u})=\lim\limits_{\epsilon\to 0}\frac{\mkern 1.5mu\overline{\mkern-1.5muC\mkern-1.5mu}\mkern 1.5mu(\boldsymbol{u}+\epsilon\delta\boldsymbol{u})-\mkern 1.5mu\overline{\mkern-1.5muC\mkern-1.5mu}\mkern 1.5mu(\boldsymbol{u})}{\epsilon},$
where $\delta\boldsymbol{u}(t)$ is an arbitrary variation defined over $t_{0},t_{f}]$. The first variation can be obtained by isolating the first order terms around the optimal solution $\boldsymbol{u}^{*}(t)$:
$\mkern 1.5mu\overline{\mkern-1.5muC\mkern-1.5mu}\mkern 1.5mu(\boldsymbol{u}^{*}+\epsilon\delta\boldsymbol{u})-\mkern 1.5mu\overline{\mkern-1.5muC\mkern-1.5mu}\mkern 1.5mu(\boldsymbol{u}^{*})\approx\delta\mkern 1.5mu\overline{\mkern-1.5muC\mkern-1.5mu}\mkern 1.5mu\big{|}_{\boldsymbol{u}^{*}(t)}(\delta\boldsymbol{u})\epsilon,$
where higher-order terms are neglected as $\epsilon\to 0$ . Since by definition $\boldsymbol{u}^{*}(t)$ is a minimum for $\mkern 1.5mu\overline{\mkern-1.5muC\mkern-1.5mu}\mkern 1.5mu$, its first variation must be zero. We first consider the deterministic case for a realization $\boldsymbol{\theta}$, where $\boldsymbol{u}_{\boldsymbol{\theta}}^{*}(t)$ is optimal for $C_{\boldsymbol{\theta}}(\boldsymbol{u})$. A well-known result in optimal control theory states that the first variation is given as [3, Eq. (5.1.13)]:
$\delta C_{\boldsymbol{\theta}}\big{|}_{\boldsymbol{u}_{\boldsymbol{\theta}}^{*}(t)}(\delta\boldsymbol{u})=\int^{t_{f}}_{t_{0}}\bigl{[}$ $\nabla_{\boldsymbol{u}}\mathcal{H}_{\boldsymbol{\theta}}^{*}\bigr{]}\delta\boldsymbol{u}(t)\operatorname{d}\!{t},$
where $\mathcal{H}_{\boldsymbol{\theta}}^{*}\coloneqq\mathcal{H}_{\boldsymbol{\theta}}\bigl{(}\boldsymbol{x}_{\boldsymbol{\theta}}^{*}(t),\boldsymbol{\lambda}_{\boldsymbol{\theta}}^{*}(t),\boldsymbol{u}_{\boldsymbol{\theta}}^{*}(t),t\bigr{)}$.
By the fundamental lemma of calculus of variations, it follows that:
$\nabla_{\boldsymbol{u}}\mathcal{H}_{\boldsymbol{\theta}}^{*}=0,\quad\text{for }t\in[t_{0},t_{f}].$
Using this result, we now turn to the mean cost and interchange the derivative and expectation:
$\delta\mkern 1.5mu\overline{\mkern-1.5muC\mkern-1.5mu}\mkern 1.5mu(\boldsymbol{u})=\delta\mathbb{E}_{\boldsymbol{\theta}}[C_{\boldsymbol{\theta}}(\boldsymbol{u})]=\mathbb{E}_{\boldsymbol{\theta}}[\delta C_{\boldsymbol{\theta}}(\boldsymbol{u})].$
@@
$=\mathbb{E}_{\boldsymbol{\theta}}\bigl{[}\int^{t_{f}}_{t_{0}}\nabla_{\boldsymbol{u}}\mathcal{H}_{\boldsymbol{\theta}}\bigl{(}\boldsymbol{x}_{\boldsymbol{\theta}}(t),\boldsymbol{\lambda}_{\boldsymbol{\theta}}(t),\boldsymbol{u}^{*}(t),t\bigr{)}\delta\boldsymbol{u}(t)\operatorname{d}\!{t}\bigr{]}$
$=\int^{t_{f}}_{t_{0}}\nabla_{\boldsymbol{u}}\mathbb{E}_{\boldsymbol{\theta}}\bigl{[}\mathcal{H}_{\boldsymbol{\theta}}\bigl{(}\boldsymbol{x}_{\boldsymbol{\theta}}(t),\boldsymbol{\lambda}_{\boldsymbol{\theta}}(t),\boldsymbol{u}^{*}(t),t\bigr{)}\bigr{]}\delta\boldsymbol{u}(t)\operatorname{d}\!{t}$
@@
By the fundamental lemma of the calculus of variations, it follows that:
@@
This shows that the optimal control $\boldsymbol{u}^{*}(t)$ is only required to satisfy the optimality conditions in the expectation of $p(\boldsymbol{\theta}\mid\mathcal{D})$, extending the minimum principle to probabilistic dynamics models that capture epistemic uncertainty.
Figure 2: Illustration of the mean Hamiltonian minimization principle. At a given time $t$, each of the $N$ samples from a probabilistic dynamics model (indexed by $\boldsymbol{\theta}$) defines a Hamiltonian $\mathcal{H}_{\boldsymbol{\theta}}\coloneq\mathcal{H}_{\boldsymbol{\theta}}\left(\boldsymbol{x}_{\boldsymbol{\theta}}(t),\boldsymbol{\lambda}_{\boldsymbol{\theta}}(t),\boldsymbol{u},t\right)$ as a function of the control $\boldsymbol{u}$ (dashed lines). The optimal control $\boldsymbol{u}^{*}$ minimizes the mean Hamiltonian, $\mathbb{E}_{\boldsymbol{\theta}}[\mathcal{H}_{\boldsymbol{\theta}}]$ (solid line), providing a control action that is optimal on average with respect to the posterior distribution over system dynamics.
@@ IV-A Trajectory sampling
[…]
$\boldsymbol{u}^{*}(t)\approx\operatorname*{argmin}_{\boldsymbol{u}}\frac{1}{M}\sum_{i=1}^{M}\mathcal{H}_{\boldsymbol{\theta}_{i}}\left(\boldsymbol{x}_{\boldsymbol{\theta}_{i}}(t),\boldsymbol{\lambda}_{\boldsymbol{\theta}_{i}}(t),\boldsymbol{u},t\right),$ (IV.1)
where $\boldsymbol{\theta}_{i}\sim p(\boldsymbol{\theta}\mid\mathcal{D})$ for $i=1,\dots,M$, and each conditioned Hamiltonian is computed along the trajectories $\boldsymbol{x}_{\boldsymbol{\theta}_{i}}(t)$ and $\boldsymbol{\lambda}_{\boldsymbol{\theta}_{i}}(t)$, which comply with the PMP conditions. An alternative formulation is to consider the mean of the optimal controls $\mkern 1.5mu\overline{\mkern-1.5mu\boldsymbol{u}\mkern-1.5mu}\mkern 1.5mu^{*}(t)=\mathbb{E}_{\boldsymbol{\theta}}[\boldsymbol{u}_{\boldsymbol{\theta}}^{*}(t)]$:
$\mkern 1.5mu\overline{\mkern-1.5mu\boldsymbol{u}\mkern-1.5mu}\mkern 1.5mu^{*}(t)\approx\frac{1}{M}\sum^{M}_{i=1}\boldsymbol{u}_{\boldsymbol{\theta}_{i}}^{*}(t).$ (IV.2)
By decoupling the optimization process over different realizations from the dynamics posterior, the mean optimal control approximation is only a valid necessary optimality condition when the Hamiltonian function is convex. However, the solutions can be computed in parallel, allowing for computationally efficient implementation since both approaches leverage Monte Carlo estimates of the state and co-state trajectories to efficiently evaluate the $M$ samples using GPU parallelization.